# Condensing and Extracting Against Online Adversaries

### v2

@@ 4 Existence of Condensers for All Values of $\ell,n$
Previous condensers from [CGR_seedless_condensers] could only show that condensers exist for uniform oNOSF sources when $\ell=o(\log n)$. They relied on existence of low-error two source extractors equipped with an additional “regularity” property. Our constructions are much simpler, recover all their results with even better parameters, and work for all values of $n$ and $\ell$, including the hardest case of $n=O(1)$.
@@ 4.2 Condenser for Two Uniform oNOSF Sources
In this subsection, we will prove Lemma 4.5. To construct the claimed condenser, we will use the following folklore result regarding existence of excellent seeded condensers (e.g., see Corollary 3 of [GLZ_cg_condenser]).
@@
We will also use the following result from [CGR_seedless_condensers] that states an adversary can’t make things too bad if it controls very few bits. We note that similar lemmas have been useful in previous construction of condensers [ben-aroya_two-source_2019, ball_randomness_2022, GLZ_cg_condenser]:
Lemma 4.9 (Lemma 6.18 in [CGR_seedless_condensers]).
Let $\mathbf{X}\sim\{0,1\}^{n}$ be an arbitrary flat distribution and let $\mathsf{Cond}:\{0,1\}^{n}\to\{0,1\}^{m}$ be such that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq k$. Let $G\subset[n]$ with $|G|=n-b$ be arbitrary. Let $\mathbf{X}_{G}\sim\{0,1\}^{n-b}$ be the projection of $\mathbf{X}$ onto $G$. Let $\mathbf{X}^{\prime}\sim\{0,1\}^{n}$ be the distribution where the output bits defined by $G$ equal $\mathbf{X}_{G}$ and remaining $b$ bits are deterministic functions of the $n-b$ bits defined by $G$ under the restriction that $\operatorname{Supp}(\mathbf{X}^{\prime})\subset\operatorname{Supp}(\mathbf{X})$. Then, $H_{\infty}^{\varepsilon^{\prime}}(\mathsf{Cond}(\mathbf{X}^{\prime}))\geq k-b$ where $\varepsilon^{\prime}=\varepsilon\cdot 2^{b}$.
@@
5 Explicit Condensers for Uniform oNOSF Sources
In this section, we will prove the following main result regarding condensing from uniform oNOSF sources , matching the existential condenser parameters of [CGR_seedless_condensers]. We state this for constant $\ell$ but our condenser can handle any $\ell=o(\log n)$.
Theorem 5.1 (Clean version of Theorem 5.5).
For constant $g,\ell$ where $g>\ell/2$, and all $n,\varepsilon$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(g,\ell\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-O(\log(m/\varepsilon))$ where $m=\Omega(n)$.
Using the transformation of low-entropy oNOSF sources to uniform oNOSF sources from Corollary 6.4 by setting $d=\ell$ (for this parameter regime, such a transformation can also be obtained using results from [CGR_seedless_condensers]), we get an explicit condenser for low-entropy oNOSF sources :
Corollary 5.2.
For constant $g,\ell$ where $g>\ell/2+1$, and all $n,k,\varepsilon$ with $k\geq\operatorname{poly}(\log(n))+\log(n/\varepsilon)+O(1)$, $\varepsilon\geq n^{-\Omega(1)}$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any ${\left(g,\ell,n,k\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))=m-O(\log(m/\varepsilon))$ where $m=k^{\Omega(1)}$.
We now use a reduction from Lemma 5.12 in [CGR_seedless_condensers]. The reduction shows how to use condensers for $g>\ell/2$ to construct condensers for all $g,\ell$ that condense up to rate $1/\left\lfloor\ell/g\right\rfloor$. [^7] The statement of the lemma in [CGR_seedless_condensers] does not explicitly state it as a reduction but such a reduction easily follows from the proof of the lemma. Using this, we construct explicit condensers for all $(g,\ell)$. Particularly, we get the following:
Corollary 5.3.
For any constant $g$ and $\ell$ and all $n,\varepsilon$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any uniform ${\left(g,\ell\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq 1/\left\lfloor\ell/g\right\rfloor\cdot m-O(\log(m/\varepsilon))$ where $m=\Omega(n)$.
We again use the transformation of low-entropy oNOSF sources to uniform oNOSF sources (Corollary 6.4) to construct a condenser for low-entropy oNOSF sources for all $g$ and $\ell$:
Corollary 5.4.
For all constant $g,\ell$, and all $n,k,\varepsilon$ with $k\geq\operatorname{poly}(\log n)$, $\varepsilon\geq n^{-\Omega(1)}$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ such that for any ${\left(g,\ell,n,k\right)}$-oNOSF source $\mathbf{X}$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))=1/\left\lfloor(\ell-1)/(g-1)\right\rfloor\cdot m-O(\log(m/\varepsilon))$ where $m=\Omega(k^{-\Omega(1)})$.
5.1 Proving the Main Theorem
Here we prove the following full version of Theorem 5.1:
Theorem 5.5.
There exists a universal constant $C$ such that for all $n,g,\ell,\varepsilon$ where $g>\ell/2$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ satisfying: for any uniform ${\left(g,\ell\right)}$-oNOSF source $\mathbf{X}$ with $g>\ell/2$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-(3C)^{\ell}\log(2\ell n/\varepsilon)$ where $m=\frac{1}{3}\cdot\left((n/2)-(3C)^{\ell}\log(2\ell n/\varepsilon)\right)$.
The proof of this theorem follows by explicitly constructing an extractor for multiple independent uniform sources, where few of them are set to constant, with the error depending on the longest source that is non-constant. Formally, we use the following:

### v3

@@ 7 Existence of Condensers for All Values of $\ell,n$
Previous condensers from [CGR24] could only show that condensers exist for uniform oNOSF sources when $\ell=o(\log n)$. They relied on existence of low-error two source extractors equipped with an additional “regularity” property. Our constructions are much simpler, recover all their results with even better parameters, and work for all values of $n$ and $\ell$, including the hardest case of $n=O(1)$.
@@ 7.2 Condenser for Two Uniform oNOSF Sources
In this subsection, we will prove Lemma 7.5. To construct the claimed condenser, we will use the following folklore result regarding existence of excellent seeded condensers (e.g., see Corollary 3 of [GLZ24]).
@@
We will also use the following result from [CGR24] that states an adversary can’t make things too bad if it controls very few bits. We note that similar lemmas have been useful in previous construction of condensers [BCDT19, BGM22, GLZ24]:
Lemma 7.9 (Lemma 6.18 in [CGR24]).
Let $\mathbf{X}\sim\{0,1\}^{n}$ be an arbitrary flat distribution and let $\mathsf{Cond}:\{0,1\}^{n}\to\{0,1\}^{m}$ be such that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq k$. Let $G\subset[n]$ with |G|=nb|G|=n-b be arbitrary. Let $\mathbf{X}_{G}\sim\{0,1\}^{n-b}$ be the projection of $\mathbf{X}$ onto $G$. Let $\mathbf{X}^{\prime}\sim\{0,1\}^{n}$ be the distribution where the output bits defined by $G$ equal $\mathbf{X}_{G}$ and remaining $b$ bits are deterministic functions of the nbn-b bits defined by $G$ under the restriction that $\operatorname{Supp}(\mathbf{X}^{\prime})\subset\operatorname{Supp}(\mathbf{X})$. Then, $H_{\infty}^{\varepsilon^{\prime}}(\mathsf{Cond}(\mathbf{X}^{\prime}))\geq k-b$ where $\varepsilon^{\prime}=\varepsilon\cdot 2^{b}$.
@@ 8 Extractors for oNOSF and oNOBF Sources via Leader Election Protocols
In this section, we provide a generic way to transform leader election and coin flipping protocols into extractors for oNOSF sources and oNOBF sources . To do so, we must formally define the online influence of coalitions.
Definition 8.1.
For any function $f:\Sigma^{\ell}\to\{0,1\}$, and any $B\subset[{\ell}]$, where $B=\{i_{1}<i_{2}<\ldots<i_{k}\}$, define $\operatorname{\mathbf{oI}}_{B}(f)$ as follows: an online adversary $\mathcal{A}$ samples a distribution $\mathbf{X}$ in online manner. It starts by sampling the variables $x_{1},x_{2},\ldots,x_{i_{1}-1}$ independently and uniformly from $\Sigma$, then picking the value of $x_{i_{1}}$ depending on $x_{<i_{1}}$. Next, the variables $x_{i_{1}+1},\ldots,x_{i_{2}-1}$ are sampled independently and uniformly from $\Sigma$, and $\mathcal{A}$ sets the value of $x_{i_{2}}$ based on all variables set so far, and so on. Define the advantage of $\mathcal{A}$ to be $\textrm{adv}_{f,B}(\mathcal{A})=|\operatorname*{\mathbb{E}}[f(\mathbf{X})]-\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell})]|$. Then $\operatorname{\mathbf{oI}}_{B}(f)$ is defined to be $\max_{\mathcal{A}}\{\textrm{adv}_{f,B}(\mathcal{A})\}$, where the maximum is taken over all online adversaries $\mathcal{A}$ that control the bits in $B$.
Now we return to our transformation from leader election and coin flipping protocols into extractors for oNOSF sources and oNOBF sources . Conceptually, given a leader election protocol, we can use an oNOSF source to simulate the protocol and then have the elected leader output its last block. We formalize this below.